# Center for Advanced Intelligence Project, RIKEN 103-0027, Tokyo, Japan

### v1

@@ 6.1 Enumerating solutions
(1.0000012030, 0.9999997855), (1.0000000063, 1.0000000010),
(1.0000000000, 1.0000000000), (0.9999998072, 1.0000000130),
which are different from each other, but very close to the exact solution $(1,1)$.
@@
    $A=\big{\{}\big{(}v_{1}(\tau),\dots,v_{n}(\tau)\big{)}:\tau\text{ is a real roots of }w\text{ with precision }\gamma_{1}\big{\}}.$
For each exact solution $\alpha$, there can be $2^{n}$ points in $A$, denoted by the set $A_{\alpha}$, which are close to $\alpha$.
@@
Going back to Example 1, we take $\gamma_{1}=10^{-6}$ and $\gamma_{2}=10^{-4}$, after Phase 2, the four approximate solutions are reduced to $(1,1)$.
@@ 6.2 Experiments on small-scale problems with worst-case scenarios
The procedure is implemented in Maple, and the results are obtained on an Intel i7 - 8665U CPU (1.90GHz) with 32 GB of RAM. The package $\mathtt{Groebner}$ can handle symbolic tasks in Algorithm 1. To compute the radical of an ideal, we use the command $\mathtt{Radical}$; to compute the reduced Gröbner basis of an ideal with respect to a given lexicographic order, we use the command $\mathtt{Basis}$. To enumerate solutions, we use numerical computations by isolating real roots of $w$. We set the variable $\texttt{Digits}:=10+\deg w$, where Digits controls the number of digits that Maple uses when making calculations with software floating-point numbers, and the second accuracy $\gamma_{2}$ is fixed $10^{-4}$.
@@ 7 Solving copositive PCPs
This section is devoted to solving another class of PCPs that may not be $\operatorname{\mathcal{D}}_{0}$, i.e., the solution set may be infinite. In particular, we focus on PCPs where involved mappings are copositive and approximate them by $\operatorname{\mathcal{D}}_{0}$ problems.
One says that a polynomial mapping $g:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ is copositive if $\left\langle\alpha\mid g(\alpha)\right\rangle\geq 0$ for all $\alpha$ in $\operatorname{\mathbb{R}}_{+}^{n}$. Assume that $\deg g=d$. Denote by $g^{\infty}$ the leading term of $g$, meaning that $g_{i}^{\infty}$ is the homogenous term of degree $d$ of $g_{i}$, for $i\in[n]$.
@@
The proof of the first statement is analogous to the proof of Proposition 2 shown in the appendix. The second one is obtained by applying directly (gowda2016polynomial, , Corollary 7.2) and (hieu2020solution, , Corollary 4.1) to polynomial complementarity problems. ∎
It is noteworthy that Proposition 4 enables us to solve approximately a PCP satisfying the assumptions therein. We take $a\in\operatorname{\mathbb{R}}^{n}$ randomly and $\|a\|$ small enough, and then solve $\operatorname{CP}(f+a)$ by using the method proposed in Section 6.1. A solution of $\operatorname{CP}(f+a)$ can be considered as an approximate solution to the original problem.
@@
We observer that $\operatorname{CP}(f+a)$, where $a=(a_{1},a_{2})$, is $\operatorname{\mathcal{D}}_{0}$ if any only if $a_{1}\neq a_{2}$. Taking $a_{1}=1/10^{6}$ and $a_{1}=1/10^{5}$ and computing solutions to $\operatorname{CP}(f+a)$ by using the proposed method, the perturbed problem has unique solution $\alpha_{a}=(0.999999,0.000000)$ which is very close to a solution $(1,0)$ of $\operatorname{CP}(f)$.
Appendix
The notions of differentiable maps and regular values can be found in monographs on differential geometry, for example, loring2011introduction .
Proof of Proposition 2. Let $\ell$ be a subset of $[n]$ and $\mathcal{V}_{\ell}$ be the set of all $\alpha\in\operatorname{\mathbb{C}}^{n}$ such that $\alpha_{i}\neq 0$ if $i\in\ell$ and $\alpha_{i}=0$ if $i\notin\ell$. Clearly, $\mathcal{V}_{\emptyset}=\{0\}$ and $\operatorname{\mathbb{C}}^{n}$ is the disjoint union of $\mathcal{V}_{\ell}$ for $\ell\subset[n]$.
Suppose that $\ell\neq\emptyset$ and consider the differentiable map $\Phi_{\ell}:\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{(\rho-1)\times n}\times\mathcal{V}_{\ell}\to\operatorname{\mathbb{C}}^{\#\ell}$ defined by
$\Phi_{\ell}(a,Q,x_{\ell})=\left(a_{i}x_{i}+Q_{i}X^{T}\right)_{i\in\ell},$
where $Q\in\operatorname{\mathbb{R}}^{(\rho-1)\times n}$ and $X$ is given in (4.1). Obviously, its Jacobian matrix with respect to $(a,Q,x_{\ell})$, denoted by $J_{\ell}$, has size $(n\cdot\rho\cdot\#\ell)\times\#\ell$. Moreover, $J_{\ell}$ has a diagonal submatrix of size $\#\ell\times\#\ell$, and the entries on the diagonal are the coordinates of $x_{\ell}$. Since all coordinates of $x_{\ell}$ are non-zero, we conclude that $J_{\ell}$ is full rank. It follows that the origin 00 in $\operatorname{\mathbb{C}}^{\#\ell}$ is a regular value of $\Phi_{\ell}$. By the Sard Theorem with parameter (dang2016well, , Theorem 2.4), there exists a generic set $\mathcal{O}_{\ell}$ in $\operatorname{\mathbb{R}}^{n\times\rho}$ such that for every $(a,Q)$ in $\mathcal{O}_{\ell}$, 00 in $\operatorname{\mathbb{C}}^{\#\ell}$ is a regular value of $\Phi_{\ell,a,Q}:\mathcal{V}_{\ell}\to\operatorname{\mathbb{C}}^{\#\ell}$ defined by $\Phi_{\ell,a,Q}(x)=\Phi_{\ell}(a,Q,x)$. It is clear that
$\Phi_{\ell,a,Q}^{-1}(0)=\{x_{\ell}:a_{i}x_{i}+Q_{i}X^{T}=0,i\in\ell\},$
and if this set is non-empty then it has finitely many points (loring2011introduction, , Theorem 9.9).
Furthermore, we observe that
$V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\subseteq\{0\}\cup\big{(}\cup_{\ell\subseteq[n],\ell\neq\emptyset}\Phi_{\ell,a,Q}^{-1}(0)\big{)},$

### v3

@@ 6.1 Enumerating Solutions
(0.9999999996, 0.9999999996), (1.0000000000, 1.0000000000),
(1.0000000000, 1.0000000000), (1.0000000100, 1.0000000100),
which are different, but very close to the actual solution $(1,1)$.
@@
    $A=\big{\{}\big{(}v_{1}(\tau),\dots,v_{n}(\tau)\big{)}:\tau\text{ is a real root of }w\text{ with precision }\gamma_{1}\big{\}}.$
For each actual solution $\alpha$, there can be $2^{n}$ points in $A$, denoted by the set $A_{\alpha}$, which are close to $\alpha$.
@@
Going back to Example 1, we take $\gamma_{1}=10^{-10}$ and $\gamma_{2}=10^{-6}$, after Phase 2, the four approximate solutions are reduced to $(1,1)$.
@@ 6.2 Experiments on Small-scale Problems with Worst-case Scenarios
The procedure is implemented in Maple, and the results are obtained on an Intel i7 - 8665U CPU (1.90GHz) with 32 GB of RAM. The package $\mathtt{Groebner}$ can handle symbolic tasks in Algorithm 1. To compute the radical of an ideal, we use the command $\mathtt{Radical}$; to compute the reduced Gröbner basis of an ideal with respect to a given lexicographic order, we use the command $\mathtt{Basis}$. To enumerate solutions, we use numerical computations by isolating real roots of $w$. We set the variable $\texttt{Digits}:=10+\deg w$, where Digits controls the number of digits that Maple uses when making calculations with software floating-point numbers, and the second accuracy $\gamma_{2}$ is fixed to $10^{-4}$.
@@ 7 Solving Copositive PCPs
This section is devoted to solving another class of PCPs that may not be $\operatorname{\mathcal{D}}_{0}$. In particular, we focus on PCPs where involved mappings are copositive and approximate them by $\operatorname{\mathcal{D}}_{0}$ problems.
One says that a polynomial mapping $g:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ is copositive if $\left\langle\alpha\mid g(\alpha)\right\rangle\geq 0$ for all $\alpha$ in $\operatorname{\mathbb{R}}_{+}^{n}$. Assume that $\deg g=d$. Denote by $g^{\infty}$ the “leading” term of $g$, meaning that $g_{i}^{\infty}$ is the homogeneous term of degree $d$ of $g_{i}$, for $i\in[n]$.
@@
The proof of the first statement is analogous to the proof of Proposition 2. The second one is obtained by applying directly (gowda2016polynomial, , Corollary 7.2) and (hieu2020solution, , Corollary 4.1) to polynomial complementarity problems. ∎
It is noteworthy that Proposition 4 enables us to solve approximately a PCP satisfying the assumptions therein: We take $a\in\operatorname{\mathbb{R}}^{n}$ randomly (then $f+a$ is $\operatorname{\mathcal{D}}_{0}$) and $\|a\|$ small enough, and then solve $\operatorname{CP}(f+a)$ by using the method proposed in Section 6.1. A solution of $\operatorname{CP}(f+a)$ can be considered as an approximate solution to the original problem.
@@
We observer that $\operatorname{CP}(f+a)$, where $a=(a_{1},a_{2})$, is $\operatorname{\mathcal{D}}_{0}$ if any only if $a_{1}\neq a_{2}$. Taking $a_{1}=1/10^{6}$ and $a_{1}=1/10^{5}$ and computing solutions to $\operatorname{CP}(f+a)$ by using the proposed method, the perturbed problem has unique solution $\alpha_{a}=(0.999999,0.000000)$ which is very close to the solution $(1,0)$ of $\operatorname{CP}(f)$.
Conclusions
We introduced a class of generic polynomial complementarity problems, named $\operatorname{\mathcal{D}}_{0}$, and proposed algorithms to compute univariate representations of their solutions. We also presented a method for solving copositive problems by approximating them with $\operatorname{\mathcal{D}}_{0}$ problems. The experimental results showed that these algorithms and method are efficient for small-scale problems.
Appendix: Proof of Theorem 4.1
The proof of Theorem 4.1 is analogous to the proof of (pham2016genericity, , Theorem 1.10), which is presented in the semi-algebraic setting. Therefore, only the most important modifications are shown below.
Proof
We only need to prove that there exists a set $\mathcal{O}$ in $\operatorname{\mathbb{R}}^{m}$ such that the three following conditions hold: $\operatorname{\mathbb{R}}^{m}\setminus\mathcal{O}$ has Lebesgue measure zero; $\mathcal{O}$ is semi-algebraic; and, for any $c\in\mathcal{O}$, 00 is a regular value of the mapping $\Phi_{c}$.
We assume without loss of generality that $\Phi^{-1}(0)$ is non-empty. It follows from the Implicit Function Theorem that $\Phi^{-1}(0)$ is a smooth sub-manifold in $\operatorname{\mathbb{R}}^{m}\times V$, where $\Phi^{-1}(0)$ is considered as a smooth manifold in $\operatorname{\mathbb{R}}^{m}\times\operatorname{\mathbb{R}}^{2n}$.
Consider the restriction of $\pi$, where $\pi:\operatorname{\mathbb{R}}^{m}\times V\to\operatorname{\mathbb{R}}^{m},(c,x)\mapsto c$ is a natural projection, on $\Phi^{-1}(0)$. By Sard’s theorem (lee2003smooth, , Theorem 6.10) (which applies to smooth mappings between real manifolds), the set of critical values of the restriction mapping $\pi|_{\Phi^{-1}(0)}$, denoted by $\mathcal{C}$, has Lebesgue measure zero in $\operatorname{\mathbb{R}}^{m}$. It is not hard to see that $\mathcal{C}$ is a semi-algebraic set in $\operatorname{\mathbb{R}}^{m}$. We define $\mathcal{O}:=\operatorname{\mathbb{R}}^{m}\setminus\mathcal{C}$. Consequently, $\mathcal{O}$ is a semi-algebraic set and $\operatorname{\mathbb{R}}^{m}\setminus\mathcal{O}$ has Lebesgue measure zero in $\operatorname{\mathbb{R}}^{m}$.
By repeating the argument in the proof of (pham2016genericity, , Theorem 1.10), we can show that 00 is a regular value of the mapping $\Phi_{c}$ for any $c\in\mathcal{O}$. ∎
Acknowledgements.
The authors would like to thank the editor and anonymous referees for their comments and João M. Pereira for his feedback on Remark 5.
Funding This work was supported by the JSPS KAKENHI Grant Number 23H03351.